Circular No. 6292

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996B IN NGC 4357 A. Dimai, Cortina d’Ampezzo, Italy, reports the discovery by Alessan-

dro Gabrijelcic (Agordo, Italy) of a possible supernova near an H II region in NGC 4357, 25′′ west and 5′′ south of the galaxy’s core. The object is of magnitude 17.5–18.0 and has been recorded on three CCD images since its discovery on Jan. 16.9, most recently on Jan. 19.9. Comparison by G. V. Williams, Harvard-Smithsonian Center for Astrophysics (CfA), of the Jan. 19.9 image with the Digital Sky Survey indeed shows a new object on the former at α = 12^{h}23^{m}56^{s}.6, δ = +48^{o}46′42′′ (equinox 2000.0), the position of the galaxy’s core having end figures 58^{s}.8, 46′′. A. Riess and R. Kirshner, CfA, also confirm the object on a V -band CCD image obtained on Jan. 21.49 UT by A. Milone with the 1.2-m reflector at Mt. Hopkins.

L. Wang, University of Texas at Austin, communicates: “A low-resolu-tion (0.5 nm) CCD spectrum (range 512–788 nm) obtained on Jan. 22 shows that this is a type II supernova with broad Hα emission and clear P-Cyg absorption. The FWHM of the Hα emission is 16 000 km/s.”

V705 CASSIOPEIAE S. P. S. Eyres, Keele University; R. J. Davis, University of Manchester,

Jodrell Bank; and M. F. Bode and H. M. Lloyd, Liverpool John Moores University, report the detection and imaging of radio emission at 6 cm from V705 Cas (Nova Cas 1993) using the MERLIN array on 1995 July 22. The source position was measured with respect to the VLBI point source 2351+550. Preliminary mapping shows the source to be extended with a peak intensity in a 50-milliarcsec beam of 370 ± 65 μJy at α = 23^{h}41^{m}47^{s}.235, δ = +57^{o}31′00′′.930 (equinox J2000.0, uncertainty ±0′′.005). The integrated flux of the source is 2140 ± 20 μJy. The angular extent of the partial shell (∼ 200 milliarcsec) indicates a distance of 2.5 kpc.

COMET P/1996 A1 (JEDICKE) Orbital elements on MPEC 1996-B04 show that this comet (cf. IAUC

6288) is of short period, with e ∼ 0.40 and P ∼ 18 years. A CCD observation by H. Mikuž (Ljubljana, Slovenia) on Jan. 17.89 UT (0.36-m reflector + V filter) gives V = 16.8.

1996 January 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden